OpenStudy (anonymous):

STATISTICS – HELP! Hypothesis Testing: ‘There is no difference between the products’ effectiveness’. 1)Using the variable Score A – Score B, conduct a two-tailed Z-test at the 5% significance level to determine if there is really a statistically significant difference between the effectiveness of each mouthwash. Check that the conditions are met to use the z-test. 2) Calculate the 95% confidence interval for the difference between the scores.

OpenStudy (anonymous):

I'm having trouble with the confidence interval part...not sure which proportion to use... see table below, regarding "Eight of the volunteers were given mouthwash A first and then mouthwash B second. The other eight volunteers were given mouthwash B first and then mouthwash A second."

OpenStudy (anonymous):

\[p^*-1.96\sqrt{\frac{ p^*(1-p^*) }{ n }} \le p \le p^*+1.96\sqrt{\frac{ p^*(1-p^*) }{ n }}\]

OpenStudy (anonymous):

I know this formula, but im not sure what p* should be...

OpenStudy (anonymous):

any ideas?

OpenStudy (perl):

the null hypothesis is that the mean of Score A - Score B is zero

OpenStudy (perl):

and you need to do a paired data test

OpenStudy (perl):

the tricky part of the problem is that they change the order of the mouthwash

OpenStudy (anonymous):

Is this a two sampZInt?

Miracrown (miracrown):

Well, it doesn't look like there are any proportions in this problem, it's just values. So you'd use the formula for that instead.

OpenStudy (perl):

@BarbaraKara Can you post the entire question with all the english. there seems to be something missing

OpenStudy (perl):

the samples are not independent samples, since it is the same person trying the two different mouthwashs

OpenStudy (perl):

I mean each subject tries two different mouthwash

OpenStudy (perl):

so i would do a 2 sample T test pooled or paired data.